### v2

# Markov switching zero-inflated space-time multinomial models for comparing multiple infectious diseases

Abstract
Univariate zero-inflated models are increasingly being used to account for excess zeros in spatio-temporal infectious disease counts. However, the multivariate case is challenging due to the need to account for correlations across space, time and disease in both the count and zero-inflated components of the model. We are interested in comparing the transmission dynamics of several co-circulating infectious diseases across space and time, where some of the diseases can be absent for long periods. We first assume there is a baseline disease that is well-established and always present in the region. The other diseases switch between periods of presence and absence in each area through a series of coupled Markov chains, which account for long periods of disease absence, disease interactions and disease spread from neighboring areas. Since we are mainly interested in comparing the diseases, we assume the cases of the present diseases in an area jointly follow an autoregressive multinomial model. We use the multinomial model to investigate whether there are associations between certain factors, such as temperature, and differences in the transmission intensity of the diseases. Inference is performed using efficient Bayesian Markov chain Monte Carlo methods based on jointly sampling all unknown presence indicators. We apply the model to spatio-temporal counts of dengue, Zika, and chikungunya cases in Rio de Janeiro, during the first triple epidemic there.
Keywords: Bayesian inference; dengue; Zika; chikungunya; Coupled hidden Markov model
1 Introduction
In epidemiology, there has been a growing interest in analyzing multivariate counts of co-circulating infectious diseases across space and time to compare transmission dynamics (Freitas et al., 2019). In this paper, we analyze bi-weekly reported case counts of Zika, chikungunya, and dengue in the 160 neighborhoods of Rio de Janeiro, Brazil, between 2015-2016, during the first triple epidemic there. The substantive aim of our analysis is to investigate how the transmission of the three diseases differs by factors such as temperature. For instance, if Zika transmits less effectively at lower temperatures compared to dengue, it would imply that its range of spread is more limited, which is helpful information for controlling the diseases (Tesla et al., 2018). Since we are mainly interested in comparing the diseases, we condition on the total number of cases in each area and bi-week and use a multinomial distribution to model their relative allocations (Dabney and Wakefield, 2005; Dreassi, 2007; Schmidt et al., 2022).
A potential issue with applying a multinomial model to multivariate spatio-temporal infectious disease counts is that many infectious diseases, especially vector-borne diseases, can be absent for long periods of time in an area (Bartlett, 1957; Coutinho et al., 2006; Adams and Boots, 2010). This behavior is illustrated in Figure 1, which shows bi-weekly reported case counts of dengue, Zika, and chikungunya in the western Rio neighborhood of Praça Seca, between 2015-2016. While dengue cases are reported consistently throughout the two years, very few Zika cases and no chikungunya cases are reported in the first year. These patterns are typical of most neighborhoods. If a disease is absent in an area, then there should be no chance of cases being reported. However, a multinomial model will always assign a positive probability to cases being reported. Therefore, there will likely be many more zeros in the counts than can be realistically produced by a fitted multinomial model (we investigate this in more detail in Section 4) (Koslovsky, 2023). Generally, when the fitted model cannot produce enough zeros, we say that the data contains excess zeros, which is very common when modeling infectious disease counts across space and time (Arab, 2015). Also, we do not want to compare the transmission dynamics of a present disease and an absent one, as it would likely overestimate the favorability of the present disease.
Figure 1: Bi-weekly reported case counts of dengue (a), Zika (b), and chikungunya (c) in Praça Seca, a west neighborhood of Rio de Janeiro, Brazil, between 2015-2016. Summer/winter seasons are highlighted in red/light blue.
When analyzing a single disease, zero-inflated count models (Lambert, 1992) are often used to deal with excess zeros in spatio-temporal infectious disease counts (Fernandes et al., 2009; Arab, 2015; Wangdi et al., 2018). With these approaches, the presence/absence of the disease is first generated through a Bernoulli process (Fernandes et al., 2009). Then, if the disease is absent in an area, it is assumed that no cases will be reported. Conversely, if the disease is present, the number of reported cases is generated through a count process, typically negative binomial or Poisson (Aktekin and Musal, 2015). A zero that arises from the count process is often interpreted as the result of the disease going undetected, referred to as a false zero (Vergne et al., 2014). This is opposed to a true zero from the Bernoulli process, which corresponds to the actual absence of the disease. Random effects can be added to both the count and Bernoulli (zero-inflated) parts of the model to account for correlations across space and time (Torabi, 2017; Giorgi et al., 2018). For example, Hoef and Jansen (2007) included random walks to account for correlations across time and conditional autoregressive effects (Besag et al., 1991) to account for correlations across space in both parts.
While the above models are popular in the univariate case, very little attention has been given to multivariate spatio-temporal zero-inflated disease modeling in general. To the best of our knowledge, no zero-inflated multinomial models have been proposed for space-time data. Pavani and Moraga (2022) and Rotejanaprasert et al. (2021) both considered zero-inflated spatio-temporal multivariate Poisson models. They used spatial and temporal random effects to account for correlations in the data. There is a close relationship between multivariate Poisson and multinomial models for multivariate disease counts (Schmidt et al., 2022). However, we prefer the multinomial approach, as it eliminates many nuisance parameters if the interest is only in comparing the diseases; see the supplementary materials (SM) Section 1. In addition, the above multivariate Poisson proposals have some important limitations that make them unsuitable for our application, including not allowing the probability of disease presence to change over time. Vector-borne diseases are highly seasonally driven and are more likely to be present in the summer and absent in the winter (Coutinho et al., 2006). Also, if an infectious disease is present in an area, it should be more likely to be present in that area and neighboring areas in the next time (Douwes-Schultz and Schmidt, 2022). The challenge with scaling up the univariate random effect models described in the previous paragraph to the multivariate case is the need to include random effects correlated across space, time and disease (not just space and time) in both the count and zero-inflated parts of the model, which would be computationally challenging.
To account for long periods of disease absence, we will extend the zero-inflated multinomial modeling framework of Zeng et al. (2022) to accommodate spatio-temporal infectious disease data. Similar to the univariate models described above, zero-inflated multinomial models allow categories (in our case diseases) to be absent from observations. If a category is absent, its multinomial probability is fixed at 0, meaning the count for that category will be 0. The presence of the categories is usually treated as random and generated through a series of independent Bernoulli processes (Koslovsky, 2023). While a zero-inflated multinomial model would allow some diseases to be absent in certain areas, existing approaches have made two independence assumptions that are inappropriate for spatio-temporal infectious disease counts. Firstly, they have assumed that the multivariate counts are independent conditional on the present categories of each observation. Infectious disease counts usually exhibit a high amount of correlation across space and time due to disease transmission (Bauer and Wakefield, 2018). Secondly, past approaches assumed that if a category is present in one observation, it does not affect whether that category or other categories are present in other observations (independence of categorical presence). If an infectious disease is present in an area, then it is more likely to be present in the next time period, as disease presence and absence are usually persistent (Coutinho et al., 2006). Also, the presence of one disease may affect whether other diseases are present due to potential disease interactions (Paul et al., 2008). Finally, due to between-area disease spread (Grenfell et al., 2001), the status of the disease in an area will affect its presence in neighboring areas.
To address the above issues, we first choose a baseline disease that is well-established in the region and assume it is always present. This avoids all diseases being absent in an area at once, which would be impossible if the total number of cases were not 0. As we discuss in Section 2.1, this problem was not considered by Zeng et al. (2022). We then assume that all other diseases switch back and forth between periods of presence and absence in each area through a series of nonhomogeneous coupled Markov chains. That is, we allow the probabilities of disease presence to depend on whether the disease was present previously in the area; whether other diseases were present previously, to account for interactions; covariates, like temperature; and the prevalence of the disease in neighboring areas, to account for spread. Therefore, unlike Pavani and Moraga (2022) and Rotejanaprasert et al. (2021), we allow the risk of disease presence to change dynamically across space, time and disease. Also, unlike Zeng et al. (2022), the presence of the categories is not independent between observations. For instance, if a disease is present in an area, this disease, and potentially other diseases, can be more likely to be present again. For the count part of our zero-inflated model, we assume that if a disease is absent in an area, no cases will be reported. If a disease is present, we assume that the odds of a new case being from that disease relative to the baseline disease (relative odds) follows an autoregressive log-linear model. We allow the log relative odds to depend on past counts in the area and neighboring areas (Fokianos and Kedem, 2003; Tepe and Guldmann, 2020). Therefore, unlike Zeng et al. (2022), we do not assume the multivariate counts are independent conditional on the presence of the diseases, and we account for the transmission of the present diseases between neighboring areas (Bauer and Wakefield, 2018). Finally, our model can be written as an autoregressive Markov switching model, meaning Bayesian inference is very tractable through the forward filtering backward sampling algorithm (Frühwirth-Schnatter, 2006).
The rest of this paper is organized as follows. Section 2 introduces our proposed autoregressive multinomial model for comparing the transmission dynamics of multiple infectious diseases in a space-time setting. Section 2.1 extends the model to incorporate zero inflation and deal with long periods of disease absence. Section 3 details our Bayesian inferential procedure, which utilizes the forward filtering backward sampling (FFBS) algorithm of Chib (1996) for efficient inference. In Section 4, we apply our model to cases of dengue, Zika and chikungunya in Rio de Janeiro during the first triple epidemic there. We close with a general discussion in Section 5.
2 An Autoregressive Multinomial Model for Comparing Transmission Dynamics
Assume we have reported case counts for $K$ different infectious diseases across $i=1,\dots,N$ areas and $t=1,\dots,T$ time periods. Let $\bm{y}_{it}=(y_{1it},\dots,y_{Kit})^{T}$ represent the vector of case counts for all diseases in area $i$ during the time interval (t1,t](t-1,t], where $y_{kit}$ represents the reported case counts for disease $k$. We will first present our model without zero inflation and then extend to the zero-inflated case in Section 2.1. Since we are mainly interested in comparing the diseases, we condition on the total number of cases and model $\bm{y}_{it}$ using a multinomial distribution,
$\bm{y}_{it}|\text{total}_{it},\bm{y}_{t-1}\sim\text{Multinom}(\bm{\pi}_{it},\text{total}_{it}),$ (1)
where $\text{total}_{it}=\sum_{k=1}^{K}y_{kit}$ and $\bm{y}_{t-1}=(\bm{y}_{1(t-1)},\dots,\bm{y}_{N(t-1)})^{T}$ is the vector of all case counts observed at the previous time point, giving the model an autoregressive structure. In (1), $\bm{\pi}_{it}=(\pi_{1it},\dots,\pi_{Kit})^{T}$ represents the expected relative distribution of the disease counts at time $t$ in area $i$ given the previously observed counts $\bm{y}_{t-1}$. If conditions in (t1,t](t-1,t] favor some diseases over others, then the share of the favored diseases should grow relative to those less favored. Therefore, we model the odds of a new case being from disease $k$ relative to the baseline disease (relative odds) as,
$\frac{\pi_{kit}}{\pi_{1it}}$ $=\lambda_{kit}^{*}=\lambda_{kit}\frac{\mathopen{}\mathclose{{\left(y_{ki(t-1)}+1}}\right)^{\zeta_{k}}}{\mathopen{}\mathclose{{\left(y_{1i(t-1)}+1}}\right)^{\zeta_{1}}},$ (2)
for $k=2,\dots,K$, where we assume disease 1 is the baseline disease, and we add 1 to the top and bottom to avoid dividing by zero, adding such constants is common in autoregressive count models (Liboschik et al., 2017; Fritz and Kauermann, 2022). The parameters $\zeta_{k}$ and $\zeta_{1}$, assumed to be between 0 and 1, are meant to account for nonhomogeneous mixing (Wakefield et al., 2019). That is, as cases increase, there tends to be a dampening effect on transmission as individuals avoid infection and governments implement interventions. In (2), $\lambda_{kit}>0$, which we link to covariates below, is a parameter that measures the favorability of conditions for the transmission of disease $k$ relative to the baseline disease during (t1,t](t-1,t]. For instance, if $\lambda_{kit}>1$ ($\lambda_{kit}<1$) then we would expect the share of disease $k$ to grow (shrink) relative to the baseline disease in (t1,t](t-1,t], after adjusting for nonhomogeneous mixing with $\zeta_{k}$ and $\zeta_{1}$. If $\zeta_{k}\approx\zeta_{1}<<1$, like for our motivating example in Section 4, then this means that if $\lambda_{kit}>1$, the share of disease $k$ will grow initially from an equal share. However, as disease $k$ becomes more dominant, the share will be pulled back towards equality as individuals avoid infection from the dominant disease.
We link $\lambda_{kit}$ to covariates and random effects using a log-linear model,
$\log(\lambda_{kit})=\alpha_{0k}+\alpha_{0ki}+\bm{x}_{kit}^{T}\bm{\alpha}_{k}+\phi_{kit},$ (3)
where $\alpha_{0ki}\sim N(0,\sigma_{k}^{2})$ is a normal random intercept meant to account for between area differences and $\bm{x}_{kit}$ is a vector of covariates that might affect the favorability of disease $k$ relative to the baseline disease. We also considered a spatially correlated intrinsic conditional auto-regressive (ICAR) model for the intercepts, $\alpha_{0ki}\sim\text{ICAR}(\sigma_{k,\text{ICAR}}^{2})$ (Besag et al., 1991). However, for our motivating example in Section 4, while our proposed model converged, we were unable to obtain convergence for some of the comparison models when using ICAR intercepts. As the results of the normal random intercept and ICAR models were very similar, see SM Section 11, we decided to use normal random intercepts in Section 4 to ensure that all comparisons were fair. We give the results of the ICAR model applied to our motivating example in SM Section 11, and the code is available on GitHub https://github.com/Dirk-Douwes-Schultz/MS_ZIARMN_code. We note that the number of cases in neighboring areas may affect the favorability of the diseases, as individuals will mix with those in neighboring areas (Grenfell et al., 2001). Therefore, following Bauer and Wakefield (2018) and Bracher and Held (2022), we allow functions of disease cases in neighboring areas, such as the prevalence across neighboring areas, to be added to the covariate vector $\bm{x}_{kit}$ to account for the geographical spread of the disease. Also, we allow previous counts of other diseases, e.g., $\log(y_{ji(t-1)}+1)$ for $j\neq k$, to be added to $\bm{x}_{kit}$ to account for potential disease interactions (Freitas et al., 2019), see Section 4 for an example.
We model the random effects $\phi_{kit}$ using a multivariate normal distribution (Xia et al., 2013; Zeng et al., 2022),
$\bm{\phi}_{it}$ $=(\phi_{2it},\dots,\phi_{Kit})^{T}\sim\text{MVN}_{K-1}(\bm{0},\bm{\Sigma}),$ (4)
where $\bm{\Sigma}$ is a K1K-1 by K1K-1 variance-covariance matrix. These random effects are mainly meant to account for overdispersion, which is a very common issue when modeling multinomial counts (Xia et al., 2013). Since $\phi_{kit}$ is correlated across diseases, they also allow for a more flexible correlation structure between the individual disease counts compared to the standard multinomial distribution; see SM Section 2.
There is an epidemiological interpretation of the parameters in Equations (1)-(4). As we show in SM Section 1, if all diseases follow a Reed-Frost chain binomial SIR model (Abbey, 1952; Vynnycky, 2010; Bauer and Wakefield, 2018) with roughly the same serial interval, the time between successive generations of cases, then we can derive Equations (1)-(4). Under this derivation, $\lambda_{kit}$ in (2) represents the ratio of the effective reproduction number of disease $k$ and the baseline disease. The effective reproduction number is an important measure of disease transmission in epidemiology, representing the average number of new infections produced by a single infectious individual in the current population before they recover (Vynnycky, 2010). The covariate effect $\alpha_{kj}$ then represents the difference between the effect of covariate $x_{kitj}$ on the effective reproduction number of disease k and the baseline disease; see SM Section 1 for more details.
While the Reed-Frost model is commonly used (Wakefield et al., 2019), it makes assumptions that are not appropriate for many diseases (Abbey, 1952), and the model can be sensitive to underreporting and reporting delay when it comes to estimating the effective reproduction number (Bracher and Held, 2021; Quick et al., 2021). Regardless, the Reed-Frost derivation does reveal an important source of potential confounding due to differences in the susceptible populations, i.e., the population of individuals who are not immune to the disease. From the Reed-Frost derivation, see SM Equation (8), ideally, we would add $\log(\delta_{ki(t-1)}/\delta_{1i(t-1)})$ to $\log(\lambda_{kit})$ in Equation (3), where $\delta_{ki(t-1)}$ is the size of the susceptible population for disease $k$ in area $i$ at time t1t-1. This makes sense intuitively; if a disease has a larger susceptible population compared to another, then transmission for that disease will naturally be favored. The effect of any covariates correlated with the ratio of the susceptible populations could then be confounded. We conduct a sensitivity analysis to attempt to control for this for our motivating example in SM Section 9.
We will call the model defined by (1)-(4) the autoregressive multinomial (ARMN) model. As detailed in the introduction, the ARMN model is not appropriate if some of the diseases are absent in an area, as it will always assign a positive probability to cases being reported. In a simulation study in SM Section 4, we show that ignoring long periods of disease absence leads to poor coverage of the 95% credible intervals and posteriors centered far from the actual parameter values. In the next subsection, we will keep the definitions from above and extend the ARMN model to deal with long periods of disease absence by adapting the proposal of Zeng et al. (2022) to our space-time setting.
2.1 Incorporating zero-inflation
Let $S_{kit}$, for $k=1,\dots,K$, be an indicator for the presence of disease $k$, so that $S_{kit}=1$ if disease $k$ was present in area $i$ during time $t$ and $S_{kit}=0$ if disease $k$ was absent. Following traditional zero-inflated epidemiological models (Vergne et al., 2014; Douwes-Schultz and Schmidt, 2022), we will assume that if a disease is absent ($S_{kit}=0$), no cases will be reported (a true or structural 0). Note that this means if all diseases are absent at once in an area ($S_{1it}=S_{2it}\dots=S_{Kit}=0$), the total number of cases would be 0 ($\text{total}_{it}=0$). However, a multinomial model conditions on the total count and does not generate it. Therefore, as it makes sense in many epidemiological applications, we will assume that the baseline disease is well established in the region and always present. That is, we assume $S_{1it}=1$ for all $i$ and $t$. We then assume that if a disease is present, the odds of a new case being from that disease relative to the baseline disease is given by Equation (2) above, leading to the following zero-inflated multinomial model,
[…]
where $\bm{S}_{it}=(S_{2it},\dots,S_{Kit})^{T}$. Note from (5) that we can have zero-reported cases of a disease ($y_{kit}=0$) even if the disease is present ($S_{kit}=1$). Following Vergne et al. (2014) and Douwes-Schultz and Schmidt (2022), we interpret this as the disease being present but undetected by the surveillance system (a false or at-risk 0).
We can compare (5) with the zero-inflated multinomial framework of Zeng et al. (2022), who assumed
$\begin{split}\pi_{1it}=\frac{S_{1it}}{S_{1it}+\sum_{j=2}^{K}S_{jit}\lambda_{jit}^{*}},\hskip 28.45274pt\pi_{kit}=\frac{S_{kit}\lambda_{kit}^{*}}{S_{1it}+\sum_{j=2}^{K}S_{jit}\lambda_{jit}^{*}},\end{split}$ (6)
for $k=2,\dots,K$. An immediate issue with (6) is that we cannot have $S_{1it}=\dots=S_{Kit}=0$ as we would divide by 0. Even if we were to define $\bm{\pi}_{it}=\bm{0}$ if $S_{1it}=\dots=S_{Kit}=0$, we would have the same issue that this implies $\text{total}_{it}=0$. There is nothing in the model of Zeng et al. (2022) to prevent $S_{1it}=\dots=S_{Kit}=0$. They may have overlooked this as they dealt with the case of $K\geq 50$ and, therefore, it is very unlikely $S_{1it}=\dots=S_{Kit}=0$. However, in our motivating example, we have $K=3$. Note that Zeng’s model (6) assuming $S_{1it}=1$ for all $i$ and $t$ is equivalent to (5).
The presence of a disease in an area should depend on whether it was present previously, as disease presence and absence are often highly persistent (Coutinho et al., 2006; Douwes-Schultz and Schmidt, 2022), and may also depend on whether other diseases were present previously, as the diseases can interact with one another (Paul et al., 2008; Sherlock et al., 2013). Additionally, the presence of a disease may depend on environmental or socioeconomic factors, such as temperature and water supply in the case of vector-borne diseases (Schmidt et al., 2011; Xu et al., 2017). Therefore, we model $S_{kit}$ as,
$\begin{split}S_{kit}|\bm{S}_{i(t-1)},\bm{y}_{t-1}&\sim\text{Bern}(p_{kit})\\[5.0pt] \text{logit}(p_{kit})&=\eta_{0k}+\bm{z}_{kit}^{T}\bm{\eta}_{k}+\underbrace{\rho_{k}^{AR}S_{ki(t-1)}}_{\begin{subarray}{c}\text{autoregressive}\\ \text{effect}\end{subarray}}+\underbrace{\sum_{j>1,\,j\neq k}\rho_{jk}^{DI}S_{ji(t-1)}}_{\begin{subarray}{c}\text{disease interaction}\\ \text{effect}\end{subarray}},\end{split}$ (7)
where $\bm{z}_{kit}$ is a vector of space-time covariates that might affect the presence of disease $k$. As the disease may spread from neighboring areas (Grenfell et al., 2001), we allow functions of disease cases in neighboring areas to be added to $\bm{z}_{kit}$, hence the dependence on $\bm{y}_{t-1}$ in (7). For instance, the prevalence of the disease in neighboring areas could be added, see Section 4. Note, as $\rho_{k}^{AR}$ in Equation (7) increases and $\eta_{0k}$ decreases, the probability of getting a consecutive period of disease presence or absence approaches 1. Therefore, (7) can account for long consecutive periods of disease presence and absence, which are often observed for infectious diseases (see Figure 1 and Coutinho et al. (2006)).
We will call the ARMN model modified by Equations (5) and (7) the Markov switching zero-inflated autoregressive multinomial (MS-ZIARMN) model. Note, from Equation (5), $\pi_{1it}+\dots+\pi_{Kit}=1$ for any possible values of the parameters and presence indicators. Also note that if diseases $k$ and $j$ are present in area $i$ during time $t$ we have $\pi_{kit}/\pi_{jit}=\lambda_{kit}^{*}/\lambda_{jit}^{*}$, which is the same as in the ARMN model described in Equations (2)-(4). That is, the MS-ZIARMN model preserves the relative distribution of the present diseases from the ARMN model. Finally, if $K=2$ and we assume Equation (5), where $\log(\lambda_{2it}^{*})=\text{logit}(\pi_{2it})=\alpha_{02i}+\bm{x}_{2it}^{T}\bm{\alpha}_{2}$ and $\text{logit}(p_{2it})=\eta_{02}+\bm{z}_{2it}^{T}\bm{\eta}_{2}$, we get the well known and popular zero-inflated binomial (ZIB) model of Hall (2000). Therefore, our model without space-time effects could be seen as a multinomial extension of Hall (2000).
3 Inferential Procedure
Here, we assume the number of diseases $K$ is small enough for matrix multiplication with a $2^{K}$ by $2^{K}$ matrix to be computationally feasible. This should be the case in most epidemiological applications; for instance, in our motivating example, we compare $K=3$ diseases. Our inferential strategy revolves around reparametrizing the MS-ZIARMN model as a Markov switching model (Frühwirth-Schnatter, 2006), which allows for very efficient Bayesian inference using the forward filtering backward sampling (FFBS) algorithm (Chib, 1996). To simplify the explanations, we will assume, without loss of generality, that $K=3$.
In the case of $K=3$, we have $\bm{S}_{it}=(S_{2it},S_{3it})^{T}$. Let $S_{it}^{*}$ be an indicator for the possible values of $\bm{S}_{it}$, so that $S_{it}^{*}=1$ if $\bm{S}_{it}=(1,1)^{T}$, $S_{it}^{*}=2$ if $\bm{S}_{it}=(0,1)^{T}$, $S_{it}^{*}=3$ if $\bm{S}_{it}=(1,0)^{T}$ and $S_{it}^{*}=4$ if $\bm{S}_{it}=(0,0)^{T}$. Then, from Equation (5),
[…]
meaning, for $K=3$, the MS-ZIARMN model is a mixture of 4 different multinomial distributions where the multinomial probabilities are fixed at 0 for the diseases absent.
Note that, from Equation (7), as $\bm{S}_{it}$ only depends on $\bm{S}_{i(t-1)}$, $\bm{y}_{t-1}$ and covariates, $S_{it}^{*}$ follows a first order nonhomogeneous Markov chain. Let $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$ be the transition matrix of $S_{it}^{*}$, where $\Gamma(S_{it}^{*}|\bm{y}_{t-1})_{jk}=P(S_{it}^{*}=k|S_{i(t-1)}^{*}=j,\bm{y}_{t-1})$ for $j,k=1,2,3,4$, $i=1,\dots,N$ and $t=2,\dots,T$. As $S_{it}^{*}$ is an indicator for $\bm{S}_{it}$, the transition matrix can be derived through
$P(\bm{S}_{it}=\bm{s}_{it}|\bm{S}_{i(t-1)}=\bm{s}_{i(t-1)},\bm{y}_{t-1})=\prod_{k=2}^{3}P(S_{kit}=s_{kit}|S_{2i(t-1)}=s_{2i(t-1)},S_{3i(t-1)}=s_{3i(t-1)},\bm{y}_{t-1}),$
for $\bm{s}_{i(t-1)},\bm{s}_{it}=(1,1),(0,1),(1,0),(0,0)$, using Equation (7). Finally, as $S_{i2}^{*}$ depends on $S_{i1}^{*}$, we require an initial distribution for the Markov chain, that is, $P(S_{i1}^{*}=j)$ for $j=1,2,3,4$ and $i=1,\dots,N$. The modeler can set the initial distribution based on how likely they believe the diseases to be present at the beginning of the study period.
Given $p(\bm{y}_{it}|\bm{y}_{t-1},S_{it}^{*},\text{total}_{it})$ and $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$ for $i=1,\dots N$ and $t=2,\dots T$, and $p(S_{i1}^{*})$ for $i=1,\dots,N$, we completely define a Markov switching model (Frühwirth-Schnatter, 2006). A Markov switching model assumes a time series can be described by several submodels, often called states or regimes, where switching between submodels is governed by a first-order Markov chain (Hamilton, 1989). In our case, the submodels are given by the 4 multinomial distributions in (8), and we switch between these submodels through the transition matrix $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$. Typically, $S_{it}^{*}$ is called the state indicator.
Let $\bm{S}^{*}=(\bm{S}_{1}^{*},\dots,\bm{S}_{T}^{*})^{T}$ be the vector of all state indicators, where $\bm{S}_{t}^{*}=(S_{1t}^{*},\dots,S_{Nt}^{*})^{T}$; $\bm{y}=(\bm{y}_{1},\dots,\bm{y}_{T})^{T}$ be the vector of all observations; and $\bm{\beta}$ the vector of all parameters in the multinomial part of the model (see (2)-(4)). Further, let $\bm{\theta}$ be the vector of all parameters in the Markov chain part of the model (see (7)), and, finally, let $\bm{v}=(\bm{\beta},\bm{\theta})^{T}$ be the vector of all model parameters. The joint distribution of $\bm{S}^{*}$ and $\bm{y}$ given $\bm{v}$ is given by,
$p(\bm{S}^{*},\bm{y}|\bm{v})=\prod_{i=1}^{N}\prod_{t=2}^{T}p(\bm{y}_{it}|\bm{y}_{t-1},S_{it}^{*},\text{total}_{it},\bm{\beta})\prod_{i=1}^{N}p(S_{i1}^{*})\prod_{t=2}^{T}p(S_{it}^{*}|S_{i(t-1)}^{*},\bm{y}_{t-1},\bm{\theta}).$ (9)
Note, from Equation (8), $S_{it}^{*}$ is only fully known when $y_{1it},y_{2it},y_{3it}>0$, in which case we must have $S_{it}^{*}=1$. It is possible, assuming $K$ is not too large, to use the forward filter (Hamilton, 1989) to completely marginalize $\bm{S}^{*}$ from (9) and calculate $p(\bm{y}|\bm{v})$, which as a function of $\bm{v}$ is the marginal likelihood function, see SM Section 3. However, we want to make inferences about $\bm{S}^{*}$ to investigate when the model believes the diseases were present or absent for model checking, as detailed in SM Section 6. Therefore, $\bm{S}^{*}$ is estimated along with $\bm{v}$ by sampling both from their joint posterior distribution $p(\bm{v},\bm{S}^{*}|\bm{y})\propto p(\bm{S}^{*},\bm{y}|\bm{v})p(\bm{v})$, where $p(\bm{v})$ is the prior distribution of $\bm{v}$. We will now discuss the prior specification in detail.
We specified a conjugate inverse-Wishart prior for $\bm{\Sigma}$ with $K$ degrees of freedom and identity matrix, as it implies the correlations in $\bm{\Sigma}$ have marginal uniform prior distributions (Gelman et al., 2013). For all non-bounded parameters, such as the intercepts and covariate effects, we used wide normal priors with a standard deviation of 5. For the precisions $1/\sigma^{2}_{k}$ we used conjugate $\text{gamma}(.1,.1)$ priors.
As the joint posterior is not available in closed form, we resorted to Markov chain Monte Carlo (MCMC) methods; in particular, we used a hybrid Gibbs sampling algorithm with some steps of the Metropolis-Hastings algorithm to sample from it. The full details of the Gibbs sampler are given in SM Section 3. We sampled all of $\bm{S}^{*}$ jointly from $p(\bm{S}^{*}|\bm{v},\bm{y})$ using the FFBS algorithm for Markov switching models (Chib, 1996). We found our MCMC algorithm mixed much faster when jointly sampling $\bm{S}^{*}$ compared to sampling each presence indicator individually. This is typically the case for Markov switching models (Frühwirth-Schnatter, 2006), and was one of our main motivations for reparametrizing the MS-ZIARMN model as a Markov switching model.
Our MCMC algorithm was implemented using the R package Nimble (de Valpine et al., 2017). We implemented the FFBS algorithm using Nimble’s custom sampler feature. All other MCMC samplers mentioned in SM Section 3 are built into Nimble. All code and data are available on GitHub https://github.com/Dirk-Douwes-Schultz/MS_ZIARMN_code. In SM Section 4, we provide a simulation study, which shows that our proposed Gibbs sampler can recover the true parameters of an MS-ZIARMN model that is specified like in our motivating example in Section 4.
4 Application to Counts of Dengue, Zika and Chikungunya
Since 2015, several countries in Latin America have experienced triple epidemics of dengue, Zika and chikungunya, placing an immense burden on the local populations (Rodriguez-Morales et al., 2016; Bisanzio et al., 2018). Between 2015 and 2016, Rio de Janeiro experienced its first triple epidemic of these diseases. Despite being transmitted by the same vector, the Aedes aegypti, Freitas et al. (2019), using space-time cluster detection methods, found that Zika and dengue clusters were much more likely in the western region of Rio (there were no chikungunya clusters found in the west). In contrast, other regions contained many clusters of each disease, though often at different points in time. However, they did not investigate how these differences in clustering might depend on covariates or other factors. Schmidt et al. (2022), using a spatial multinomial regression model, found that cases of chikungunya were more likely in urban areas and Zika cases were less likely in population-dense areas, compared to dengue cases. However, a spatial analysis cannot investigate how space-time factors might have affected differences in the transmission intensity of the diseases, which is of great interest in this example. For instance, laboratory studies have shown that the transmission of Zika may be more sensitive to temperature compared to dengue, which is important as it implies Zika will not be able to spread as effectively into North America and the highlands (Tesla et al., 2018).
Therefore, we apply our MS-ZIARMN model to cases of the three diseases, $K=3$, across the 160 neighborhoods of Rio de Janeiro, $N=160$, for each bi-week between 2015-2016, $T=52$. Our model seems appropriate for this example, as there appear to be long periods of disease absence in both the time series of chikungunya and Zika cases, see Figure 1 for example (overall 65% of Zika cases are equal to 0 and 75% of chikungunya cases are equal to 0).
4.1 Model specification and fitting
We modeled the Rio data with a bi-weekly time step as this roughly corresponds to the serial interval of the three diseases (Majumder et al., 2016; Riou et al., 2017). We took dengue as the baseline disease (k=1). Recall, from Section 2.1, that this means we are assuming that dengue was present in every neighborhood of Rio throughout the study period. We believe this is a reasonable assumption, as dengue has been circulating in Rio since 1986 and is well established there (Teixeira et al., 2009). We then considered $k=2$ to represent Zika and $k=3$ to represent chikungunya.
For $\bm{x}_{kit}$ in (3), covariates that may affect the favorability of Zika or chikungunya transmission relative to dengue transmission, we first took all spatial covariates considered in Schmidt et al. (2022): the percentage of neighborhood $i$ covered in green area, $\text{verde}_{i}$; the social development index of neighborhood $i$, $\text{SDI}_{i}$; and the population density of neighborhood $i$, $\text{popdens}_{i}$. We additionally considered the percentage of neighborhood $i$ that was occupied by favelas (slums), $\text{favela}_{i}$. These spatial covariates were chosen as the transmission of arboviral diseases can be significantly influenced by socioeconomic factors such as water supply and urbanicity (Schmidt et al., 2011), see Schmidt et al. (2022) for more details. As for space-time covariates, we first took the average weekly maximum temperature in neighborhood $i$ during bi-week $t$, $\text{temp}_{it}$. Temperature is an important factor to consider, since if the transmission of Zika or chikungunya is more sensitive (less sensitive) to temperature compared to dengue, their range of spread will be smaller (bigger) (Tesla et al., 2018; Mercier et al., 2022). Note that if there are many more cases of Zika or chikungunya in neighboring areas compared to dengue, then we would expect the share of those diseases to grow, all else being equal, due to between-area mixing (Stoddard et al., 2013). Therefore, we considered the previous prevalence of disease $k$ and disease $1$ (dengue) across neighboring areas, $\log\mathopen{}\mathclose{{\left(\sum_{j\in NE(i)}y_{kj(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}}\right)$ and $\log\mathopen{}\mathclose{{\left(\sum_{j\in NE(i)}y_{1j(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}}\right)$, in $\bm{x}_{kit}$, where $NE(i)$ is the set of neighboring areas of neighborhood $i$ and $\text{pop}_{j}$ is the population of neighborhood $j$. We considered two areas neighbors if they shared a border. To simplify the model, we assumed the effect of neighboring dengue prevalence was the same on both $\log(\pi_{2it}/\pi_{1it})$ and $\log(\pi_{3it}/\pi_{1it})$, like with the within-area autoregression. Finally, it was speculated in Freitas et al. (2019) that Zika circulation could have been inhibiting chikungunya transmission. To account for these kinds of disease interactions, we considered $\log(y_{3i(t-1)}+1)$ in $\bm{x}_{2it}$ and $\log(y_{2i(t-1)}+1)$ in $\bm{x}_{3it}$. That is, chikungunya cases were allowed to affect the favorability of Zika transmission relative to dengue transmission, and Zika cases were allowed to affect the favorability of chikungunya transmission relative to dengue transmission.
As for covariates potentially associated with the presence of Zika or chikungunya, $\bm{z}_{kit}$ in (7), we considered mostly the same factors included in $\bm{x}_{kit}$. An exception is that we did not include cases of the other diseases, as the Markov chain defined in (7) already accounts for disease interactions through $\rho_{jk}^{DI}$. Also, since the presence is not defined relatively, we did not include neighboring dengue prevalence.
We also need to specify the initial state distributions, see Section 3. As there were no chikungunya cases reported at all in the first year, we assumed chikungunya had only a 5 percent chance of being present at the start of 2015. For Zika, there were a few, but not many, cases reported through most of 2015; therefore, we assumed a 10 percent chance of initial presence. We fitted the MS-ZIARMN model specified above to the Rio data using our proposed Gibbs sampler from Section 3. We ran the Gibbs sampler for 250,000 iterations on three chains with an initial burn-in of 50,0000 iterations. All sampling was started from random values in the parameter space to avoid convergence to local modes. Convergence was checked using the Gelman-Rubin statistic (<<1.05), the minimum effective sample size (>>1000) and by visually examining the traceplots (Plummer et al., 2006). The traceplots for all parameters (except the random effects) are shown in SM Section 10. It took around a day to run three chains of the Gibbs sampler in parallel on a laptop with an i9-12900HK, 2500 MHz, CPU.
For comparison purposes, we also considered a model without zero inflation to ensure that there is justification for treating the many zeroes in the data, as hypothesized. That is, the ARMN model described in Equations (1)-(4) with the same covariates $\bm{x}_{2it}$ and $\bm{x}_{3it}$ as the MS-ZIARMN model. We also considered a zero-inflated model without the coupled Markov chains, that is, with $\rho_{k}^{AR}=0$ and $\rho_{jk}^{DI}=0$ for all $k$ and $j>1,\,j\neq k$ in Equation (7), which we will call the ZIARMN model. Without the coupled Markov chains, our model is a finite mixture and not a Markov mixture model, and the inference is greatly simplified, e.g., there is no need to use the FFBS algorithm. Finally, we compared with the zero-inflated multinomial model of Zeng et al. (2022) as our approach is built on their framework. Zeng et al. (2022) assumed that the probability a category was absent from an observation only depended on the category, so we let $\text{logit}(p_{kit})=\eta_{0k}$ for $k=2,3$ in Equation (7). We kept the rest of the model the same as the MS-ZIARMN model. Note that Zeng et al. (2022) did not consider any covariates or space-time structure in the multinomial probabilities. However, this is inappropriate for our example and did not fit the data well (results not shown). Also, Zeng et al. (2022) did not assume the baseline category was always present; however, this appears to be necessary when modeling a small number of categories, as discussed in Section 2.1.
Table 1 shows the widely applicable information criterion (WAIC) (Gelman et al., 2014) of the 4 considered models. We calculated the WAIC by marginalizing out the state indicators $\bm{S}^{*}$ as recommended by Auger-Méthé et al. (2021), see SM Section 5 for more details. Note that the model with the smallest WAIC is considered to have the best fit, and as a rule of thumb, a difference of 5 or more in the WAIC is considered significant. Table 1 shows it is not only important to account for zero inflation, but also covariates and correlations across space, time and disease in the zero-inflated process (i.e., the process generating disease presence). Indeed, the model of Zeng et al. (2022) had a worse fit compared to the ARMN model (which does not account for zero-inflation), despite the other zero-inflated models having a much better fit.
Table 1: Shows the WAIC of the 4 considered models from Section 4.1 fitted to the Rio data. The best fitting model, the one with the lowest WAIC, is bolded.
Model Zeng (2022) ARMN ZIARMN MS-ZIARMN
WAIC 23,986 23,974 22,756 22,091
4.2 Results
Table 2 shows the estimated coefficients from the multinomial part of the fitted MS-ZIARMN model, that is, the intercepts and covariate effects from Equation (3). From the intercept row, at average values of the covariates, Zika transmission was favored over dengue and chikungunya transmission. That is, under average conditions, the share of Zika cases tended to grow over time in an area (at least from an equal share, see the nuances in Section 2). It is interesting to compare these estimated intercepts with those from the ARMN and Zeng (2022) models in SM Table 1. Unlike the MS-ZIARMN model, the Zeng (2022) and ARMN models both estimated that, on average, Zika and chikungunya transmission was much less intense in Rio compared to dengue transmission. This does not correspond to our knowledge of the epidemiology of these diseases and is likely due to a failure to properly account for the many zeroes. Zika is generally considered more transmissible due to Aedes aegypti transmitting Zika at a higher rate (Freitas et al., 2019). This further illustrates it is important to account for zero inflation and to model the zero inflation properly (e.g., account for correlations).
Table 2: Posterior means and 95% posterior credible intervals (in parentheses) for the estimated coefficients from the multinomial part of the fitted MS-ZIARMN model. The intercept row shows $\lambda_{kit}$ for $k=2$ (Zika) and $k=3$ (chik.) in a typical area at average values of the covariates. Recall, if $\lambda_{kit}>1$ ($\lambda_{kit}<1$) then the share of disease $k$ relative to dengue will tend to grow (shrink) over time. Other rows show the ratio of $\pi_{kit}/\pi_{1it}$ (relative odds ratio) or the ratio of $\lambda_{kit}$ (rate ratio) (both are the same, see Equations (2)-(3)) corresponding to a unit increase in the covariate. All covariates are standardized. Significant effects are bolded. See Section 4.1 for an explanation of the covariates.
Relative Odds Ratio or Rate Ratio
Covariates Parameter Zika-dengue chik.-dengue
Intercept $\exp({\alpha_{0}})$ 1.14 (1.03, 1.26) 1.02 (.91, 1.13)
$\text{verde}_{i}$ $\exp({\alpha_{\text{verde}}})$ 1.02 (.95, 1.09) .92 (.85, 1)
$\text{SDI}_{i}$ $\exp({\alpha_{\text{SDI}}})$ 1.07 (.99, 1.15) 1.01 (.92, 1.11)
$\text{popdens}_{i}$ $\exp({\alpha_{\text{popdens}}})$ 1.02 (.94, 1.11) 1.06 (.96, 1.16)
$\text{favela}_{i}$ $\exp({\alpha_{\text{favela}}})$ .98 (.91, 1.05) .94 (.87, 1.03)
$\text{temp}_{it}$ $\exp({\alpha_{\text{temp}}})$ 1.14 (1.09, 1.20) .85 (.80, .90)
Neighborhood dengue prevalence $\exp({\alpha_{\text{NeiDeng}}})$ .70 (.67, .74) .70 (.67, .74)
Neighborhood Zika prevalence $\exp({\alpha_{\text{NeiZika}}})$ 1.59 (1.48, 1.70)
Neighborhood chik. prevalence $\exp({\alpha_{\text{NeiChik}}})$ 1.43 (1.36, 1.50)
Previous Zika cases $\exp({\alpha_{\text{Zika}}})$ .90 (.85, .96)
Previous chik. cases $\exp({\alpha_{\text{Chik}}})$ .98 (.95, 1.02)
Interestingly, unlike in Schmidt et al. (2022), we did not find strong evidence that any of the spatial covariates were important for explaining differences in the transmission of the diseases. We found that Zika transmission was favored at warmer temperatures and chikungunya transmission was favored at colder temperatures, compared to dengue transmission. Both of these findings have been observed in laboratory studies (Tesla et al., 2018; Mercier et al., 2022). Indeed, a concern with chikungunya is that it could spread into colder areas of Europe that are unreachable by dengue (Mercier et al., 2022). To better quantify the temperature effects, we plot $\lambda_{kit}$, from Equation (2), versus temperature for $k=2$ (Zika) and $k=3$ (chikungunya) in Figure 2. From the figure, we have strong evidence that, under otherwise average conditions, Zika transmission was favored above 33.2 degrees and chikungunya transmission was favored below 32 degrees, compared to dengue transmission (the average temperature in Rio during the study was 33.8 degrees, with the middle 95 percent being from 27.2 to 39.1 degrees).
Figure 2: Shows the posterior means (black solid line) and 95% posterior credible intervals (black dashed lines) of $\lambda_{kit}$ versus temperature for $k=2$ (Zika) (left) and $k=3$ (chikungunya) (right). We assumed an average area, other covariates fixed at their average values and a null space-time random effect. Recall, if $\lambda_{kit}>1$ ($\lambda_{kit}<1$) then the share of disease $k$ relative to dengue will tend to grow (shrink) over time. The red horizontal dashed lines are drawn at temperatures below and above which $\lambda_{kit}$ is significantly different from 1.
The neighborhood prevalence effects are by far the strongest in Table 2 for explaining differences in the relative distributions of the diseases. For instance, we estimated that a standard deviation increase in the log of the prevalence of Zika across neighboring areas during the previous week was associated with a 59 (48, 70) percent increase in the odds of a new case being Zika relative to dengue. As Aedes aegypti primarily bite during the day, its arboviruses can spread rapidly between areas (Stoddard et al., 2013). We found that areas with more Zika cases were less favorable towards the transmission of chikungunya relative to dengue. That is, we estimated that a standard deviation increase in the log of the previous number of Zika cases in an area was associated with a 10 (4, 15) percent decrease in the odds of a new case being chikungunya relative to dengue. This does not necessarily mean that Zika transmission inhibited chikungunya transmission in a non-relative sense. It could be that Zika transmission encouraged both dengue and chikungunya transmission, but the effect on dengue transmission was larger. Therefore, this result needs to be interpreted with some caution.
We estimated $\zeta_{1}$, $\zeta_{2}$ and $\zeta_{3}$ from Equation (2) as .44 (.38, .49), .36 (.31, .32) and .43 (.36, .51), respectively, indicating a high degree of nonhomogeneous mixing. That is, the model predicts that if one of the diseases becomes very dominant in a neighborhood, the relative distribution of the diseases will move towards a uniform distribution, perhaps because individuals avoid infection from the dominant disease. However, as avoiding one of the diseases should help avoid the others, as they are transmitted by the same vector, $\zeta_{k}$ could also be mainly reflecting differences in disease immunity, which are difficult to account for explicitly (see SM Section 9). From Equations (3)-(4), we estimated the standard deviation of $\phi_{2it}$ as .75 (.71, .79), of $\phi_{3it}$ as .80 (.75, .86) and the correlation between $\phi_{2it}$ and $\phi_{3it}$ as .59 (.52, .65). This implies that the correlations between the disease cases, marginalizing the random effects and conditional on their total, is about what we would expect without the random effects, see the simulation study in SM Section 2. That is, the random effects do not induce any excess correlations between the disease counts. However, they do induce a large amount of overdispersion, as seen in Figure 4.
Table 3: Posterior means and 95% posterior credible intervals (in parentheses) for the estimated parameters from the Markov chain part of the fitted MS-ZIARMN model. The intercept row shows the probability of Zika presence and the logit of the probability of chikungunya presence, assuming no previous disease presence in the area, no cases in neighboring areas and at average values of the other covariates. The Zika column shows the ratio of $p_{2it}/(1-p_{2it})$ (odds of Zika presence) corresponding to a unit increase in the covariate. The chik. column shows the untransformed estimates (not exponentiated, see text). All covariates are standardized except the neighborhood prevalences, which are given in units of 1 case per 25,000 (average combined size of neighboring areas) at small prevalences.
Odds Ratio Untransformed
Covariate Parameter Zika presence chik. presence
$\text{Zika col.}\,\,\,|\,\,\,\text{chik. col.}$
Intercept $\exp(\eta_{0})/(1+\exp(\eta_{0}))\,\,\,|\,\,\,\eta_{0}$ .1 (.08, .12) -9.06 (-11.97, -7)
$\text{verde}_{i}$ $\exp(\eta_{\text{verde}})\,\,\,|\,\,\,\eta_{\text{verde}}$ 1.14 (.98, 1.31) 08 (-.20 .37)
$\text{SDI}_{i}$ $\exp(\eta_{\text{SDI}})\,\,\,|\,\,\,\eta_{\text{SDI}}$ 1.11 (.96, 1.29) .12 (-.18, .42)
$\text{popdens}_{i}$ $\exp(\eta_{\text{popdens}})\,\,\,|\,\,\,\eta_{\text{popdens}}$ 1 (.85, 1.19) .02 (-.33, .37)
$\text{favela}_{i}$ $\exp(\eta_{\text{favela}})\,\,\,|\,\,\,\eta_{\text{favela}}$ 1.03 (.89, 1.19) 0 (-.3, .29)
$\text{temp}_{it}$ $\exp(\eta_{\text{temp}})\,\,\,|\,\,\,\eta_{\text{temp}}$ 1.44 (1.24, 1.66) 1.11 (.6, 1.7)
Neighborhood Zika prevalence $\exp(\eta_{\text{NeiZika}})\,\,\,|\,\,\,\eta_{\text{NeiZika}}$ 1.70 (1.55, 1.87)
Neighborhood chik. prevalence $\exp(\eta_{\text{NeiChik}})\,\,\,|\,\,\,\eta_{\text{NeiChik}}$ 1.09 (.85, 1.35)
Previous Zika presence $\exp(\rho_{2}^{AR})\,\,\,|\,\,\,\rho_{23}^{DI}$ 6.41 (3.95, 9.94) 5.12 (3.2, 8)
Previous chik. presence $\exp(\rho_{32}^{DI})\,\,\,|\,\,\,\rho_{3}^{AR}$ 3.46 (2.42, 4.98) 11.55 (8.13, 16.12)
Table 3 shows the estimated parameters from the Markov chain component of the model, that is, from Equation (7). The Zika column shows the ratio of the odds of Zika presence corresponding to a unit increase in the covariate. However, the intercept for chikungunya is highly negative, making the odds ratios difficult to interpret. Therefore, we did not transform the chikungunya estimates. From the intercept row, it was very unlikely for chikungunya to emerge in an area under average conditions. Indeed, chikungunya emergence was only likely after Zika had emerged and temperatures rose in the next summer. In contrast, the emergence of Zika was much more random, occurring with a 10 (8, 12) percent chance bi-weekly under average conditions. Once chikungunya did emerge, it was likely to stay in the area (probability of .88 (.63, .99) assuming no neighboring cases, no Zika presence in the previous bi-week and otherwise average conditions). In contrast, Zika persistence was more reliant on the presence of cases in neighboring areas (there was only a 41 (32, 49) percent chance of Zika being present if it was present previously and there were no neighboring cases). Both Zika and chikungunya were more likely to be present at high temperatures and when there were cases in neighboring areas. The effects of neighborhood prevalence are especially large, again illustrating the diseases’ high rate of geographical spread. The presence of Zika was an indicator of chikungunya presence, and vice versa. This could be unrelated to disease interactions and due to unmeasured confounding, as the diseases share the same vector.
Some of the chikungunya estimates in Table 3 are extreme. Chikungunya cases were first reported in most neighborhoods between bi-weeks 28 and 34 (measured from the beginning of 2015). This implies a significant change in the probability of chikungunya presence around bi-week 28 across all neighborhoods. Therefore, the extreme estimates in Table 3 are likely due to the model attempting to fit this large and fast shift in the probability of chikungunya presence. In light of the above, the chikungunya estimates in Table 3 should be interpreted with caution, as there may be unmeasured confounders driving the change in the risk of chikungunya presence around bi-week 28. For example, spread from travel outside the city (Nunes et al., 2015), which is difficult to measure and may be related to factors such as temperature. In contrast, Zika cases emerged much more randomly across the city, as reflected in Table 3.
Figure 3: Shows the posterior mean of $\bar{\lambda}_{ki}=\mathopen{}\mathclose{{\left(1/\sum_{t=2}^{T}S_{kit}}}\right)\sum_{t=2}^{T}\lambda_{kit}S_{kit}$} for $k=2$ (Zika) (top map) and $k=3$ (chik.) (bottom map). If $\bar{\lambda}_{ki}>1$ ($\bar{\lambda}_{ki}<1$) then the transmission of disease $k$ was favored (disfavored) on average over dengue when it was present in the neighborhood, and the share of disease $k$ tended to grow (shrink) relative to dengue in the neighborhood.
The maps in Figure 3 show the posterior mean of $\bar{\lambda}_{ki}=\mathopen{}\mathclose{{\left(1/\sum_{t=2}^{T}S_{kit}}}\right)\sum_{t=2}^{T}\lambda_{kit}S_{kit}$} (average value of $\lambda_{kit}$ when disease $k$ was present in neighborhood $i$) for $k=2$ (Zika) (top map) and $k=3$ (chik.) (bottom map). From the bottom map, chikungunya transmission was disfavored in the west of the city when it was present, and the share of chikungunya cases tended to decline there. This was partly due to temperatures being 2-3 degrees higher on average in the west (see SM Figure 2) and to chikungunya transmission being disfavored at warmer temperatures relative to the other two diseases (see Table 2). Also, there were many Zika cases in the west of the city, and chikungunya transmission was disfavored in areas with many Zika cases. Zika transmission was favored throughout most of the city when it was present. This can be partly explained by the inherent favorability of Zika, i.e., the intercept in Table 2; by temperatures typically being higher in Rio when Zika was present (the temperature covariate was centered using the overall average temperature); and by Zika transmission being favored at higher temperatures.
Figure 4: Top graphs show the posterior means (solid lines) and 95% posterior credible intervals (shaded areas) of the fitted values $y_{kit}^{\prime}$, see SM Section 6. The bottom graphs show the posterior probability that the disease was present in each bi-week, $P(S_{kit}=1|\bm{y})$ versus $t$. "MS-ZIARMN no random effect" refers to the MS-ZIARMN model fit without including $\phi_{kit}$ from Equations (3)-(4).
Figure 4 shows fitted values in the top graphs of each panel from three models. SM Section 6 explains how we constructed the fitted values by drawing new space-time random effects and simulating from the fitted model. The bottom graphs of each panel in Figure 4 show the posterior probability the disease was present in the area, $P(S_{kit}=1|\bm{y})$; see SM Section 6. If an area reports cases of disease $k$, then we know that it is present there, that is, $P(S_{kit}=1|\bm{y})=1$ if $y_{kit}>0$. If an area does not report cases, the disease could be absent or present but undetected, the probability of the latter is given by $P(S_{kit}=1|\bm{y})$. By comparing panels (a) and (b), the random effects $\phi_{kit}$, from Equations (3)-(4), are an important component of the model and induce a large amount of necessary overdispersion. From panel (c), the model of Zeng et al. (2022), which does not account for correlations in disease presence, estimates chikungunya as always present in Catumbi, except for the initial state (this is also true for Zika and in all other areas). Therefore, the model presented in Zeng et al. (2022) is too restrictive to capture the likely complex space-time distribution of disease presence in the city. Panel (d) shows Zika likely circulated undetected at the end of 2016. One way to think of $P(S_{kit}=1|\bm{y})$ (from the bottom graphs) is as a weight for the observation’s contribution to the likelihood of the multinomial parameters of the model. For example, from panel (a), the MS-ZIARMN model does not use any chikungunya observations before around week 28 in Catumbi to inform the multinomial estimates in Table 2. In contrast, the model of Zeng et al. (2022), from panel (c), will use all observations to inform the multinomial model, which likely causes bias due to the large number of 0s, see the discussion at the beginning of Section 4.2 and our simulation study in SM Section 4.
The model takes disease absence to mean there was no chance of cases being reported. However, if the diseases are present in very small amounts, the chance of cases being reported could be close enough to 0 that it is indistinguishable. Indeed, if we replaced the 0s in Equation (8) by say .005, and normalized the probabilities, we would likely get similar parameter estimates. Therefore, disease presence should be interpreted closer to "a feasible chance of cases being reported" than "the presence of at least one infection".
5 Discussion
We have proposed a zero-inflated spatio-temporal multinomial model for comparing the transmission dynamics of multiple co-circulating infectious diseases. Our approach can account for long periods of disease absence while investigating how certain factors are related to differences in the transmission intensity of the diseases. Past zero-inflated multinomial models (Tang and Chen, 2019; Zeng et al., 2022; Koslovsky, 2023) have made independence assumptions that are inappropriate for spatio-temporal infectious disease counts, such as assuming categorical presence was independent between observations. We accounted for correlations across space, time and disease in both the multinomial and zero-inflated components through a combination of autoregression, regressing on past observations, and by assuming the presence of the diseases followed a series of coupled Markov chains. This approach allowed for efficient and computationally feasible Bayesian inference using the FFBS algorithm. We also showed, in a simulation study in SM Section 4, that failing to account for zero-inflation often leads to many important parameters not being covered by the credible intervals, which further lends support to our approach.
We assume that the cases of the diseases present in an area jointly follow a multinomial model, conditioning on their total. It is more standard in the disease mapping literature to jointly model multiple disease counts with some form of a multivariate Poisson distribution (Jack et al., 2019). There are advantages to both approaches and we give a more detailed comparison of the two (without model fitting) in SM Section 1.2. If one is only interested in comparing the transmission of the diseases, like here, then a multinomial model eliminates many nuisance parameters and any shared space-time factors. A multivariate Poisson approach could estimate the overall effects of a covariate on transmission, not just the relative effects, and does not need to assume one of the diseases is always present. As mentioned in the introduction, only Pavani and Moraga (2022) and Rotejanaprasert et al. (2021) have considered zero-inflated spatio-temporal multivariate Poisson models. Both relied on random effects to capture correlations in the data. However, these proposals have some important limitations, including not allowing the probabilities of disease presence to change over time. These modeling assumptions are inappropriate for our motivating example and likely for others as well. For instance, Figure 4 clearly shows the probabilities of disease presence changing over time. An approach based on random effects could become intractable as one would need to consider random effects correlated across three dimensions, space, time and disease, in both the count and zero-inflated components of the model. Our approach seems much more computationally feasible and could be easily extended to the multivariate Poisson case, see SM Section 1.2. We will focus on these extensions and comparisons with the multinomial model in future work.
In our application, we found that Zika generally had more intense transmission compared to dengue and chikungunya, but was also not able to transmit as well at a lower temperature. This was not a major factor in tropical Rio, see Figure 3, but could be important in North America and Europe. In contrast, chikungunya transmission was relatively higher at lower temperatures, meaning it might fare better in colder regions compared to Zika and dengue. Laboratory studies have come to similar conclusions (Tesla et al., 2018; Mercier et al., 2022). Alternative models that did not account for zero inflation or did not model correlations in disease presence (Zeng et al., 2022) did not fit as well and produced estimates inconsistent with our knowledge about the epidemiology of the diseases.
There are also some important limitations of our work. Being an observational study, unmeasured confounding could be an important issue, especially concerning differences in the susceptible populations, which are unobservable. We tried using differences in cumulative incidence as a proxy for differences in the susceptible populations, in a sensitivity analysis in SM Section 9, and our results did not change. However, differences in cumulative incidence could be an imperfect proxy due to, for example, changes in reporting rates across space and time. Another limitation is that weather covariates, such as temperature, can have lagged and non-linear effects on the transmission of arboviral diseases (Lowe et al., 2018). As our multinomial model estimates differences in effects on transmission, our approach might be somewhat robust to non-linearity if the non-linear patterns are similar between diseases. It would be challenging to incorporate non-linear and lagged effects into our already complex modeling scheme. Shrinkage methods could help deal with the many parameters that would need to be added to the model by eliminating unimportant effects (Wang et al., 2023). Also, the effect of weather covariates on the incidence of arboviral diseases may change across space or time (Nobre et al., 2005). To accommodate this, the covariate effects could be modeled as random walks (Nobre et al., 2005) or other spatio-temporal random effects in future work. However, this would be computationally challenging since covariates can enter the model in each of the relative odds and each of the probabilities of presence. Finally, another limitation of our approach is that a multinomial model is not appropriate for predicting future disease counts. To predict future cases of a disease, a multinomial model would need to know the total number of cases across all diseases in the future. A possible solution is to build a separate model for the total count (Schmidt et al., 2022), or a multivariate Poisson model could be used, as discussed above.
Acknowledgements
This work is part of the PhD thesis of D. Douwes-Schultz under the supervision of A. M. Schmidt in the Graduate Program of Biostatistics at McGill University, Canada. Douwes-Schultz is grateful for financial support from IVADO and the Canada First Research Excellence Fund/Apogée (PhD Excellence Scholarship 2021-9070375349). Schmidt is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999) and from Institut de valorisation des données (IVADO) (PRF-2019-6839748021). L. P. Freitas is supported by the Visiting Researcher Program of the Sergio Arouca National School of Public Health (ENSP), Oswaldo Cruz Foundation (Fiocruz). M. S. Carvalho acknowledges support from the National Council for Scientific and Technological Development (CNPq) - Brazil [307450/2021-0] and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) - Brazil [E-26/203.967/2024]. This research was enabled in part by support provided by Calcul Québec (www.calculquebec.ca) and Compute Canada (www.computecanada.ca).